A roll of wire 105cm long wraps exactly three times around a triangular piece of wood. If the triangle has sides with lengths 2x cm, 5x-3 and 3x+8cm, find the dimensions of the triangle.
Answer by checkley77(12844)

3*2x+3(5x-3)+3(3x+8)=105
6x+15x-9+9x+24=105
30x=105+9-24
30x=90
x=90/30
x=3
2*3=6cm. for one side.
5*3-3=15-3=12cm. for the second side.
3*3+8=9+8=17cm. for the third side.
proof:
3*6+3*12+3*17=105
18+36+51=105
105=105